If one-half the smaller number of two numbers is increased by one-third the larger number, the result is 2. If one-third the smaller number increased by one-ninth the larger number, the result is 1. Find the numbers.
in worksheet!!

Answer by edjones(8007)

Let s=smaller, b=bigger
s/2 + b/3 = 2
s/3 + b/9 = 1
.
3s+2b=12 multiply each side by 6 to eliminate fractions.
3s+b=9 multiply each side by 9 to eliminate fractions.
--------subtract bottom from top
==>b=3
3s+2*3=12 substitute 3 for b in the 1st equation.
3s+6=12
3s=6
s=2
Check:
2/2+3/3=2
2/3+3/9=1
.
Ed